# Water-rich sub-Neptunes and rocky super Earths around different Stars: Radii shaped by Volatile Partitioning, Formation, and Evolution

### v1

@@ 4.1 Key quantities for comparison of mass-radius relations
To better compare the models and observational data, we introduce here four quantities which are not expected to be influenced by observational biases and should relate to different, stellar-mass-dependent processes shaping the period-mass-radius distribution of small planets.
The upper-mass end of the rocky planet distribution . To not rely on outliers, this is measured as the 90th quantile of the rocky planet mass distribution. Rocky planets are identified by having a normalized density of at least 0.65 times that of an equal-mass Earth composition planet (using the relation by Zeng et al., 2016).
@@
The volatile-planet mass-radius relation slope . This can be fitted using least-squares fitting as implemented in the scipy python package (Virtanen et al., 2020). By using a single slope, we assume that the mass-radius relation is of the form . To allow for an estimation of the error, we use a bootstrapping approach for the observed data where for 1000 repetitions, an equal-sized distribution of planets is drawn with replacing from the observed data. In addition, we randomize the values of the observed data by sampling from a skewed normal distribution based on the errors in mass and radius space. For synthetic data, we obtain comparable error estimates by drawing 1000 times a sample of planets with the same number of points as the observed data (using a Jackknife approach, i.e. without replacing).
@@
Fig. 10 provides a means for comparison of mass-radius distributions with corresponding power-law fits to the volatile distribution as well as visualizing the quantities and . We find that all models predict a smaller value than the observed , match within errors , and predict a significantly shallower slope of the volatile planet mass-radius relation than the observed . We find that the quantity is in our model shaped by mass-loss and not imprinted from formation. The population of water-rich planets after formation extends to lower masses (i.e. to approximately 1 M⊕ smaller than ). Since the values lie below the observed ones, this hints at too inefficient photoevaporation caused by the prescription itself or by too high mean molecular weight atmospheres in the modeled planets.
While it can be a useful probe for photoevaporation, the values for could be influenced by the observational bias. In contrast, the slope of the mass-radius relation is a less bias-dependent indication that too many planets with high metallicity envelopes are present in the models. This can also be seen from a bifurcation in the volatile-rich population with a group clustered at relatively low radii and another with larger radii. The lower population is made from almost pure steam worlds which is at odds with observation. Although, some planets populate this region (e.g. GJ 9827 d, Passegger et al., 2024), they are not as commonly observed as predicted by the model. This group is even more abundant in the Water Sequestration and Layered models. We note that the analysis for the Water Sequestration model is made more difficult as many water-rich planets with small outgassed atmospheres are classified to be part of the rocky population.
The qualitative observation that is captured by all models. If all rocky planets originate from being stripped from an equal-composition massive envelope, there should be an irradiation trend visible with rocky planets close to orbiting on short orbital periods and volatile-rich planets at the low-end of the volatile-rich mass-radius distribution an wide orbits. We visually inspect the data for such a trend and do not find any correlation in the observed data. In the synthetic data from the Mixed model, however, the most massive rocky planets indeed orbit on short orbital periods and there are more long-orbit volatile-rich planets around .
@@
We can further try to identify trends with stellar masses for the mass-radius characteristics. In Fig. 11, the observational data and Mixed and Water Sequestration model output is split into two stellar mass bins with a separation at 0.6 M⊙. One apparent effect is that the two characteristic masses and lie closer to each other around lower stellar masses for the synthetic and observed data. Instead of an overlap in mass space of the volatile-rich and rocky population, they show a clear transition at 4 M⊕. In the Mixed model, this is mainly because the rocky planet population decreases in mass. It is to a large degree sourced by planets which formed interior to the iceline and this population shrinks with decreasing stellar mass (Burn et al., 2021). The findings here seem to contradict the theoretical findings of the work by Venturini et al. (2024). They found for their pebble-accretion based planets (with structures analogous to the Mixed model) that the overlap of the volatile-rich planets decreases with increasing stellar mass – opposite to what we find. However, this difference is because we applied a cut at the runaway greenhouse limit to our synthetic planets (Sect. 2.3). The planets with radii larger than that expected from condensed water are also not present around low mass stars in the work of Venturini et al. (2024, their Fig. 1). Indeed, when excluding planets cold enough to allow for condensed water, their data reveals a similar transition from rocky to volatile-rich planets at 3-4 M⊕ around low-mass stars.
The comparison suggests that our Mixed model could be improved if planets have lower initial envelope metallicities: the slope of the volatile-rich planet mass-radius relation would steepen and the would increase towards a generally better match with observations. The solution is however not to dissolve large quantities of water in the magma ocean and metallic core as this water will be later outgassed and remain on too low-mass planets in disagreement with observations. Instead, it is more likely that less water or more H/He should be accreted during formation. The latter option is unlikely as the formation model already includes optimistic estimates on the envelope cooling efficiency (via reduced assumed opacities, Mordasini, 2014). Potential mechanisms for a reduction of water accretion could be radioactive heating and subsequent loss of water from planetesimals (Lichtenberg et al., 2019) or recycling of water back to the ambient gas (Moldenhauer et al., 2022; Wang et al., 2023). However, with a reduction of the atmospheric metallicity the radius valley location would shift to even larger radii (Mordasini, 2020) in disagreement with Kepler data (see Sects. 3.3 and 3.4). Thus, the process which reduces the water content might require a planet mass-dependency which does not affect low-mass planets but more higher mass ones, such as expected for recycling (Wang et al., 2023).
@@ 4.2 Negligible effect of fractionation
In this work, we included the effect of fractionating outflows of oxygen relative to hydrogen for the evolution of synthetic exoplanets (see Appendix A for the implementation). This effect was studied in the context of exoplanets and coupled to evolutionary calculations by Hu et al. (2015), Malsky & Rogers (2020); Malsky et al. (2023), Gu & Chen (2023), and Cherubim et al. (2024) although for He or deuterium. Since oxygen is heavier, the effect should be more pronounced. However, we did not find a significant effect on the radius.
@@
Since the initial envelope masses are considerably massive, following the output of our formation modeling, a large photoevaporative mass loss is required to alter their radius significantly. Such a flow is however not in the regime where fractionation occurs as can be seen from Eq. (14), where the mass loss is in the denominator of the subtrahend. Thus, for , which evaluates to a relative mass loss of , fractionation does not occur and mass of oxygen is lost proportionally to its fraction (Hunten et al., 1987; Zahnle & Kasting, 1986; Zahnle et al., 1990). While this estimate is based on analytical theory, the numerical investigation of Johnstone (2020) found a similar estimate of this so-called critical flux. This is satisfied for most planets in our simulation until the last stage of the evolution (see Sect. 3.1) in general agreement with the findings of Cherubim et al. (2024). However, in that study, He-rich atmospheres remained in some cases. Although this difference should be investigated in more details, it could be due to their study also sampling small initial envelope fractions on the percent level. This speculation is supported by the study of Gu & Chen (2023) who found a dependency on the initial gas envelope mass for deuterium fractionation. Such low envelope fractions are not obtained from our formation modeling and are therefore not included in this study but could be candidates for fractionation to occur to a more significant degree.
@@
We do not find a significant influence of the estimate of the upper atmosphere number density (see Appendix A). Further extension of the fractionation model to different species, such as He, and including a realistic temperature and flow profile should be investigated in the future (see also the discussions of deviations from analytical estimates in Johnstone, 2020; Schulik & Booth, 2023).
@@
There are observational prospects to probe the atmospheres of planets at the upper edge of the radius valley (Malsky et al., 2023; Cherubim et al., 2024). Since He and deuterium are unlikely to be outgassed from the interior, they provide a probe for the process of fractionation occurring. However, a future observational confirmation of the absence of He-dominated atmospheres should not be over-interpreted as being caused by the complete absence of photoevaporation. Instead, the planet could have had a more massive initial gaseous envelope and photoevaporation could be efficiently acting and also remove the He content of the planet. Rather, the presence of He or deuterium rich atmospheres could be used as a probe for initial gaseous envelope masses.
@@ 4.3 Outlook on models including sequestration
Here, we report comparisons to observational data which do not favor the most physically advanced Water Sequestration model. However, this finding should be put into relation to the developments required for a comprehensive model with atmosphere-interior coupling.
@@
First, the Water Sequestration model includes only the dissolution of water but excludes that of the other modeled constituents, namely He and hydrogen. For He, this could be justified, but hydrogen dissolution into the interior is expected to occur (Hirschmann et al., 2012; Kite et al., 2019). Moreover, the dissolved material is bound to undergo chemical reactions. Instead of outgassing water, the secondary envelope can be of a different composition. However, if large quantities of oxygen were accreted via water ice and correspondingly, the mantle is in a high redox state, water is indeed expected to also be the main constituent of the outgassed atmosphere (Bower et al., 2022; Lichtenberg et al., 2023). Nevertheless, given how sensitive photoevaporative mass-loss is on the radius and therefore the molecular weight of the atmosphere, more outgassed species should be included. In particular, gas accretion predicts that the hydrogen to oxygen ratio increases with increasing planetary mass (Mordasini et al., 2016), therefore, the evolution of the high-mass planets in our simulation would change more significantly and outgassed hydrogen and other species could improve the match with observations in this regime.
@@
Another neglected effect is the temperature dependency of the solubility. For simplicity, we assumed a constant exterior temperature of 3000 K for the computation of the water sequestration. Instead, the planet would cool and the retention of water in the magma ocean would therefore decrease (Chachan & Stevenson, 2018). This would imply earlier outgassing than found here and potentially more efficient loss of water from the planet. Qualitatively, this effect is not expected to change the mass-radius relation of the Water Sequestration model, it would only reduce the average radius of the closely-correlated, water-rich, (almost) envelope-free planets.
@@
A potentially more important related effect is however the solidification of the magma ocean (Hamano et al., 2013; Bower et al., 2018, 2022). Depending on the details of crystallization pattern and viscosity (e.g. whether or not a solid lid forms at the top of the mantle), most of the volatile budget could be removed from the mantle. If mantle solidification occurred on hot planets up to four Earth masses, this effect might reconcile with observations the low-mass end of the Water Sequestration model results presented here.
@@
An exploration in this direction is also motivated by the findings of Gaidos et al. (2024) who explore the age dependency of the small planet population. They suggest that the atmosphere-interior coupling could produce effects on Gyr timescales as observed and further motivate an inclusion of water condensation. Qualitatively, water condensation could have a similar effect as switching from the Mixed to the Layered scenarios and could be quantified using these approaches as a basis (see also the discussion in Venturini et al., 2024).

### v2

@@ 4.1 Key quantities for comparison of mass-radius relations
To better compare the models and observational data, we introduce here four quantities which are not expected to be heavily influenced by observational biases and should relate to different, stellar-mass-dependent processes shaping the period-mass-radius distribution of small planets.
The upper-mass end of the rocky planet distribution . To not rely on outliers, this is measured as the 90th quantile of the rocky planet mass distribution. Rocky planets are identified by having a normalized density of at least 0.65 times that of an equal-mass Earth composition planet (using the relation by Zeng2016).
@@
The volatile-planet mass-radius relation slope . This can be fitted using least-squares regression as implemented in the scipy python package (scipy). By using a single slope, we assume that the mass-radius relation is of the form . To estimate uncertainties, we use a bootstrapping approach for the observed data: for 1000 repetitions, an equal-sized distribution of planets is drawn (with replacement) from the observations. In addition, we randomize the observed values by sampling from a skewed normal distribution based on the errors in mass and radius space. For synthetic data, we obtain comparable error estimates by drawing 1000 subsamples of planets with the same number of points as the observed data, using a Jackknife approach (i.e. without replacement).
@@
Fig. 10 provides a means for comparison of mass-radius distributions with corresponding power-law fits to the volatile distribution as well as visualizing the quantities and . We find that all models predict a smaller value than the observed , match within errors , and predict a significantly shallower slope of the volatile planet mass-radius relation than the observed . We find that the quantity is in our model shaped by mass-loss and not imprinted from formation: The population of water-rich planets after formation extends to lower masses than (i.e. to approximately 1 M⊕). Since modeled lie below the observed one, this hints at too inefficient photoevaporation caused by the prescription itself or by too high mean molecular weight atmospheres in the modeled planets.
While it can be a useful probe for photoevaporation, the values for could be influenced by the observational bias. In contrast, the slope of the mass-radius relation is a less bias-dependent indication that too many planets with high metallicity envelopes are present in all models. This can also be seen from a bifurcation in the volatile-rich population with a group clustered at relatively low radii and another with larger radii. The lower population is made from almost pure steam worlds. Although, some planets populate this region (e.g. GJ 9827 d, Passegger2024), they are not as commonly observed as predicted by all models. This group is more abundant in the Water Sequestration and Layered models compared to the Mixed one.
The qualitative observation that is captured by all models. If all rocky planets originate from being stripped from an equal-composition massive envelope, there should be an irradiation trend visible with rocky planets close to orbiting on short orbital periods and, vice-versa, volatile-rich planets around on wide orbits. We visually inspect the data for such a trend and recover it for the Mixed model but do not find any correlation in the observed data.
@@
We also examine mass–radius trends with stellar mass by splitting the observational data and the Mixed and Water Sequestration model output into two bins, separated at 0.6 M⊙ (Fig. 11). One apparent effect is that the two characteristic masses and lie closer to each other around lower stellar masses for the synthetic and observed data. Instead of an overlap in mass space of the volatile-rich and rocky population, they show a clear transition at 4 M⊕. In the Mixed model, this is mainly because the rocky planet population decreases in mass. It is to a large degree sourced by planets which formed interior to the iceline and this population shrinks with decreasing stellar mass (Burn2021). The findings here seem to contradict the theoretical findings of the work by Venturini2024. They found for their pebble-accretion based planets (with structures analogous to the Mixed model) that the overlap of the volatile-rich planets decreases with increasing stellar mass – opposite to what we find. However, this difference is because we applied a cut at the runaway greenhouse limit to our synthetic planets (Sect. 2.3). The planets with radii larger than that expected from condensed water are also not present around low mass stars in the work of Venturini2024. Indeed, when excluding planets cold enough to allow for condensed water, their data reveals a similar transition from rocky to volatile-rich planets at 3-4 M⊕ around low-mass stars.
The comparison suggests that our Mixed model could be improved if planets have lower initial envelope metallicities: the slope would steepen and the would increase towards a generally better match with observations. This is partially the case for the highest masses in the Sequestration model (see Sect. 3.2.4). Other possible options are that less water or more H/He should be accreted during formation. Within the standard assumptions on planet formation, the latter option is unlikely as the formation model already includes optimistic estimates on the envelope cooling efficiency (via reduced assumed opacities, Mordasini2014) which should not lie lower and thus cannot increase the H/He content at fixed mass. However, it is possible to accrete also more H/He at fixed mass if part of the metals reside in the envelope (fully consistent with the Mixed model, see Venturini2016 and the discussion in Sect. 4.4.2).
Potential mechanisms for a reduction of water accretion could be radioactive heating and subsequent loss of water from planetesimals (Lichtenberg2019) or recycling of water back to the ambient gas (Moldenhauer2022; Wang2023). However, with a reduction of the atmospheric metallicity the radius valley location would shift to even larger radii (Mordasini2020) in disagreement with Kepler data (see Sects. 3.3 and 3.4). Thus, the process which reduces the water content might require a planet mass-dependency which does not affect low-mass planets but more higher mass ones, such as expected for recycling (Wang2023).
@@ 4.2 Negligible effect of fractionation
In this work, we included the effect of fractionating outflows of oxygen relative to hydrogen for the evolution of synthetic exoplanets (see Appendix A for the implementation). This effect was studied in the context of exoplanets and coupled to evolutionary calculations by Hu2015, Malsky2020; Malsky2023, Gu2023, and Cherubim2024 although for He or deuterium. Since oxygen is heavier, the effect should be more pronounced. However, we did not find a significant effect on the radius.
@@
Since the initial envelope masses are considerably massive, following the output of our formation modeling, a large photoevaporative mass loss is required to alter their radius significantly. Such a flow is however not in the regime where fractionation occurs as can be seen from Eq. (14), where the mass loss is in the denominator of the subtrahend. Thus, for , which evaluates to a relative mass loss of , fractionation does not occur and mass of oxygen is lost proportionally to its fraction (Hunten1987; Zahnle1986; Zahnle1990). While this estimate is based on analytical theory, the numerical investigation of Johnstone2020 found a similar estimate of this so-called critical flux. This is satisfied for most planets in our simulation until the last stage of the evolution (see Sect. 3.1) in general agreement with the findings of Cherubim2024. However, in that study, He-rich atmospheres remained in some cases. Although this difference should be investigated in more details, it could be due to their study also sampling small initial envelope fractions on the percent level. This speculation is supported by the study of Gu2023 who found a dependency on the initial gas envelope mass for deuterium fractionation. Such low envelope fractions are not obtained from our formation modeling and are therefore not included in this study but could be candidates for fractionation to occur to a more significant degree.
@@
We do not find a significant influence of the estimate of the upper atmosphere number density (see Appendix A). Further extension of the fractionation model to different species, such as He, and including a realistic temperature and flow profile should be investigated in the future (see also the discussions of deviations from analytical estimates in Johnstone2020; Schulik2023).
@@
There are observational prospects to probe the atmospheres of planets at the upper edge of the radius valley (Malsky2023; Cherubim2024). Since He and deuterium are unlikely to be outgassed from the interior, they provide a probe for the process of fractionation occurring. However, a future observational confirmation of the absence of He-dominated atmospheres should not be over-interpreted as being caused by the complete absence of photoevaporation. Instead, the planet could have had a more massive initial gaseous envelope and photoevaporation could be efficiently acting and also remove the He content of the planet. Rather, the presence of He or deuterium rich atmospheres could be used as a probe for initial gaseous envelope masses.
@@ 4.3 Outlook on models including sequestration
Here, we report comparisons to observational data which do not favor the most physically advanced Water Sequestration model, at least at planetary masses below 3 M⊕. However, this finding should be put into relation to the developments required for a comprehensive model with atmosphere-interior coupling.
@@
First, the Water Sequestration model includes only the dissolution of water but excludes that of the other modeled constituents, namely He and hydrogen. For He, this could be justified, but hydrogen dissolution into the interior is expected to occur (Hirschmann2012; Kite2019). Moreover, the dissolved material is bound to undergo chemical reactions. Instead of outgassing water, the secondary envelope can be of a different composition (see the recent work by Werlen2025, for equilibrium chemistry results). However, if large quantities of oxygen were accreted via water ice and correspondingly, the mantle is in a high redox state, water is indeed expected to also be the main constituent of the outgassed atmosphere (Bower2022; Lichtenberg2023). Nevertheless, given how sensitive photoevaporative mass-loss is on the radius and therefore the molecular weight of the atmosphere, more outgassed species should be included. In particular, gas accretion predicts that the hydrogen to oxygen ratio increases with increasing planetary mass (Mordasini2016), therefore, the evolution of the high-mass planets in our simulation would change more significantly and outgassed hydrogen and other species could improve the match with observations in this regime.
@@
Another neglected effect is the temperature dependency of the solubility. For simplicity, we assumed a constant exterior temperature of 3000 K for the computation of the water sequestration. Instead, the planet would cool and the retention of water in the magma ocean would therefore decrease (Chachan2018). This would imply earlier outgassing than found here and potentially more efficient loss of water from the planet. Of a similar order of magnitude but with inverse effects is the simplification that pressures used for sequestration include only the water envelope mass and ignore the H/He content. A consistent pressure at the magma surface would slightly increase water fractions in the melt for planets with significant H/He content. We note that the presence of water on planets resulting from the formation model is usually either negligible or significant in mass which then leads to water being the dominant constituent in the atmosphere, and thus determining the total pressure.
@@
A potentially more important related effect is however the solidification of the magma ocean (Hamano2013; Bower2018; Bower2022). In our approach, the whole mantle material is assumed to remain in molten state. However, depending on the details of crystallization pattern and viscosity (e.g. whether or not a solid lid forms at the top of the mantle), most of the volatile budget could be removed from the mantle upon solidification. If it occurred on hot planets up to three Earth masses, this effect might in fact reconcile with observations the low-mass end of the Water Sequestration model results presented here since outgassing would decrease the bulk density of the planet leading to more efficient, likely complete, loss of volatiles and thus removal of planets from the low-density, low-mass region where no real planets are found.
@@
An exploration in this direction is also motivated by the findings of Gaidos2024 who explore the age dependency of the small planet population. They suggest that the atmosphere-interior coupling could produce effects on Gyr timescales as observed and further motivate an inclusion of water condensation. Qualitatively, water condensation could have a similar effect as switching from the Mixed to the Layered scenarios and could be quantified using these approaches as a basis (see also the discussion in Venturini2024).
4.4 Other Uncertainties and Caveats
4.4.1 Formation model parameters
When synthesizing planetary populations, several parameters are not tightly constrained. Therefore, different pathways can lead to populations with reasonable reproduction of the observed exoplanet bulk mass, total radius, and distance distributions. A review highlighting this challenge, listing parameter choices in the literature, and highlighting their effect on the planet population was recently compiled (Burn2024a). In particular, the migration prescription combined with the chosen disk initial conditions and viscosity sensitively affects whether the observable small planet population contains migrated planets or those formed in-situ.
The formation simulations extended here assume a viscous disk with standard prescriptions for type I migration from Paardekooper2011 using the same viscosity as for the disk evolution. The discussion on whether and to which degree disks can be modeled as viscous (i.e. their mechanisms driving turbulence can be modeled with an effective viscosity) is ongoing. There is a realistic possibility that inner disk regions (au) turn relatively laminar during extended stages in disk evolution. This would likely trigger earlier gap-opening of planets with the effect of reduced migration of sub-Neptune mass planets (e.g. Ogihara2018, see also discussion in Burn2024a). We note that the creation of migration prescriptions for low-viscosity disks is an active research topic and need to be included in future global simulations for more accurate conclusions (Paardekooper2023; Ziampras2024; Aoyama2023; Weder2025).
While more efficient migration, or similarly, a more shallow radial distribution of planetesimals, promotes the abundance of water-rich sub-Neptune planets, it does not have a direct impact on the pattern in mass-radius space as the growth history of the individual planets remains the same. In the analysis done here, with the exception of Fig. 9, we did not put strong emphasis on the number of sub-Neptunes compared to super Earths.
The changes in mass radius space, in our simulations, are governed by the planetary composition and the partitioning of elements discussed here. It is therefore important to review the effects on bulk composition of a planet, which is naturally related to the initial composition of the disk. For example, the maximum water abundance in solids is limited by the disk water abundance – which in turn is, in the outer disk, governed by the partitioning of the available oxygen to icy species. The observed ice abundances towards star forming regions show significant scatter (Boogert2015). Here, they were not varied between stars (the initial abundances follow Marboeuf2014b) although there could be significant intrinsic scatter. Furthermore, the stellar abundance was kept Solar and the oxygen fraction in minerals was similarly fixed (following Thiabaud2014). Moreover, we assumed that no refractory organics are present. The addition of some refractory carbon species (pure amorphous carbon or organic compounds such as PAH) is strongly favored by the observed carbon depletion in diffuse interstellar clouds (Whittet2003). Depending on the composition and ratio of carbon to oxygen in the refractory organics, the ice abundance can vary. If carbon is condensed as amorphous carbon, water abundance might be promoted, but more likely is a depletion of all ices (including water) due to condensation into refractory organics. Furthermore, the water abundance could be negatively affected by physical processes (radioactive heating, recycling) discussed above.
Finally, the results presented here are in principle sensitive to the bulk abundance of H/He to water on the planet. This ratio is sensitive to the efficiency of gas accretion which in turn depends on energy input (i.e. solid accretion rate), boundary conditions varying with the location in and model of the disk, and assumed envelope opacities. We note that here, we chose low envelope opacities during formation assuming grain growth following (Mordasini2014). However, as discussed in Mordasini2020, Venturini2024, and other locations in this manuscript, there is a balancing effect in evolutionary calculations with photoevaporation: larger H/He contents inflate the planet size which leads to more efficient mass loss and thus smaller final planet sizes. Due to these competing effects, at short orbital periods, a higher H/He content leads to smaller final planets while at larger distances (or larger planetary mass), the opposite effects holds. Another effect of more efficient H/He accretion would be that planets more easily reach runaway gas accretion and the formation of giants would be promoted, which in turn requires a re-tuning of other parameters in a population synthesis exercise to match the observed abundance of giants.
Overall, we expect the mass-radius relationship – in shape – to have relatively small sensitivity on typical planet formation parameters which keep the composition fixed but be rather sensitive to gas accretion efficiency and assumptions on solid compositions (e.g. ice to dust ratio). However, the exact abundances of planets in the mass radius distribution (e.g. number of sub-Neptunes to super Earths) is sensitive to parameters affecting migration rates and growth efficiencies such as viscosity and properties of the solid planet building blocks. Future works should disseminate these dependencies in more detail.
4.4.2 Mixing in the formation stage
In this work, we assume for the Mixed model that mixing occurs after the formation stage. For ease of implementation and computational efficiency, we did not mix H/He with water during the assembly of the planets in the formation stage. If mixing occurs, it likely emerges as soon as certain threshold conditions are met (Innes2023). A disk-embedded planet with a mixed envelope structure has different properties compared to a layered structure. Its higher density implies that the mass resides deeper in the potential well which favors gas accretion while cooling could be negatively affected by higher molecular opacities due to water vapor. If opacities are still mainly dust-dominated, that is, they are not sensitive to water opacities, a positive effect on gas accretion results (Venturini2015; Venturini2016). This implies that the H/He to water ratio in the planets modeled here could be too low and future model generations need to allow for consistent treatment of the partitioning during formation and evolution. This could help to improve the match with observations for the Mixed model as discussed in Sect. 4.
4.4.3 Collisions
Giant impacts among young planets are expected to significantly affect the planets. Depending on impact mass ratios and velocities, different outcomes can be expected, such as disruption of the bodies, merging, or no efficient mass transfer (hit-and-run) (e.g. Cambioni2019). In terms of bulk composition this leads to compositional mixing with more intermediate-composition planets emerging. This aspect is included in the simulations presented here where perfect merging of all components except H/He envelopes is assumed. However, atmospheres present are also significantly affected (Schlichting2015; Schlichting2018). For this aspect, we could not use a fully consistent treatment of collisions even though we use multiple-body simulations as basis (see Sect. 2.2.1). This has two impacts on our results.
First, atmospheric mass-loss is likely underestimated in cases where collisions occurred after the disk dissipation. While the atmosphere of the smaller collision partner is always removed from the mass budget, we assumed no effect on the mass of the atmosphere of the target. Venturini2020 discuss in detail the effect of using a model for atmospheric mass removal due to collisions and found a general shift of the population towards higher densities.
Second, for the Layered model, mixing of the atmosphere is likely to occur after a giant impact. This would imply that the planet transitions from a Layered scenario to the Mixed case. From our simulations, we expect this effect to affect the population differently depending on their atmospheric escape rates. If the timescale of mass-loss is shorter than the few Myr on which the collisions occur, H/He is already lost at the time at which the impacts occur. This is the case for the close-in, low-mass population which is completely stripped of H/He and is making up most of the observed population below M⊕. For the planets where atmospheres are Mixed by impacts, they would become more resistant to H/He loss. It is noteworthy that giant collisions are also common during the disk stage which further promotes the Mixed scenario over the Layered one.